# FAST functional connectivity implicates P300 connectivity in working memory deficits in Alzheimer’s disease

### v1

@@ III-B Application to visual short-term memory binding in Alzheimer’s disease
EEG micro-states are transient patterns of the electroencephalogram that occur in very small temporal windows and are considered to be related to the most basic of human neurological processes. They have been previously shown to be able to distinguish between neurological disorders such as schizophrenia based on these tiny temporal window differences where the overall functional connectivity of the brain may be very similar [21]. Recently, there has been significant interest in these EEG micro-states in neurological disorder diagnosis. Chu et al.[22] showed that combining graph metrics based on dynamic functional connectivity in small temporal windows with typical classification algorithms showed significantly improve performance in the early detection of Parkinson’s Disease (PD), showing the discriminating ability of these EEG micro-states.
As mentioned earlier, we have a large number of options for the choice of bi-variate node function and long-term stable connectivity filter, but it is essential to prevent data dredging. From our simulations, we found that the FAST Correlation filter with the pairwise squared difference as the node-function to be the most effective combination at detecting relevant ERP’s. Thus this will be the basis of our main results in our FAST Connectivity Analysis.
We used FAST connectivity matrices computed from participants in both VSTM tasks (shape and binding) to filter mean matrices of patients and controls in the binding VSTM task in both data sets. The temporal heat-maps of the theta band in the familial (top) and sporadic (bottom) dataset comparing controls (left) and patients (right) are shown below in Figure 4.
Figure 4: Clustering Coefficient Heatmaps of mean FAST filtered control (left) and patient (right) matrices in the binding VSTM task in the familial and sporadic AD group in the theta band
It is interesting to note that in the FAST filtered mean matrix of patients against controls we see patches of increased clustering coefficient values in patients consistently across both datasets. Given the nature of the squared difference this could indicate points of increase in anti-correlative information or variation between signal pairs.
@@
While the mean FAST filtered matrices do contain valuable information we do lose some transient information. We thus ran experiments for the Familial AD and Sporadic AD datasets separately with a single FAST filter computed from participants in both the shape and binding tasks for each frequency band. As mentioned in our statistical methods section we computed the mean clustering coefficient and edge weights of each participant at ten disjoint temporal windows of the total one second epoch of interest and undertook non-parametric statistical testing between controls and patients to look for temporal windows where there is concurrently a significant difference between controls and patients in the binding task and no significant difference in the shape task. This exploits the proposed binding deficit established in [9]. Figure 5 shows the plots of the log of the $p$-values of the patients versus controls in shape and binding tasks for the familial AD dataset with values below the black dotted line indicating a significant difference. The scale is provided for the delta plot.
Refer to caption Refer to caption
@@
Figure 5: FAST Connectivity Shape vs Binding task in delta(top),theta(middle) and alpha(bottom) bands for the Familial Group
@@
The first thing we notice is that the behaviour of the delta and theta band in the binding task is nearly identical with significant results found at 0.3-0.4 seconds by the mean edge weights and 0.4-0.5 by the mean weighted clustering coefficient. We can see how having two different network metrics can aid the detection of clinically significant results. We see the shape tasks for these time steps are not significantly different thus these can be considered results of clinical interest as the specificity of binding deficits observed behaviorally are replicated here at a neural level. The binding task in the alpha band seems to show some behaviour similar to the theta and delta bands with a clinically interesting result at 0.4-0.5 seconds, however this would not pass FDR correction. The beta band seems to follow the same pattern in the 0.3-0.6 range with a dip towards the significance line in the binding task and movement away from it in the shape task. In light of volume conduction effects, the Gamma band was found to yield spurious significant results, consequently warranting its exclusion from our analysis. Overall, we notice a trend of clinically significant results in the 0.3-0.6 second range in the familial AD data-set, mainly in the lower frequency bands.
We now perform our FAST Connectivity Analysis on the Sporadic AD data set. Recall this is a completely independent dataset with a 128 electrode setup with 17 control and 13 (older) patients at high risk of Sporadic AD. The corresponding plots of the $p$-values between patients and controls for Shape and binding tasks over the one second epoch are given below. Note there may be a slight discrepancy between interval values due to the differing sampling rates, e.g. 0.3-0.4 seconds is actually 0.29-0.39 seconds.
Refer to caption Refer to caption
@@
Figure 6: FAST Connectivity Shape vs Binding task in delta(top),theta(middle) and alpha(bottom) bands for the Sporadic Group
@@
We notice similar patterns in the lower frequency bands with the binding task in the delta and theta band following similar patterns in the mean edge weights and the network metrics picking up clinically interesting results in the theta band at 0.4-0.5 and 0.5-0.6 seconds. This is an overlap in the 0.3-0.6 second range with the familial AD data set. The delta band has a clinically interesting result at 0.5-0.6 seconds with the mean edge weights and a highly significant result at 0.8-0.9 seconds using the weighted clustering coefficient,this could be related to the emotion related Late Positive Potential (LPP). The LPP, characterized by a gradual positive shift in activation, typically manifests approximately 400-1000 ms following stimulus presentation. Its amplification has been linked to memory encoding and retention mechanisms [9]. Furthermore, it has been correlated with post-retrieval phases, such as decisional and evaluation processes which could be affected by AD.Again the overlap with the delta band in the familial AD group in the 0.3-0.6s range is seen. Moreover,the alpha band has clinically interesting results at 0.1-0.2 seconds with the weighted clustering coefficient and 0.2-0.3s and 0.5-0.6s with the mean edge weights. It seems the similarity to the theta band in the familial AD group is growing with the behaviour in the 0.3-0.6s range being much more prominent in the alpha band. The beta band has a significant result at the 0.4-0.5s time-step which is similar to the behaviour of the beta band in the familial AD data-set (dip towards significance line in binding task, movement away in the shape task). Overall, there are consistent overlapping results of clinical interest in the 0.3-0.6s temporal range.
@@
We then applied FDR correction at the 10 percent and 5 percent level to account for multiple comparisons. Table 2 below shows the time intervals at which there is a significant difference between controls and patients in the binding task and not in the shape task.
@@
TABLE II: FDR Corrected Data sets. The underlined text represents the Binding p-value below 0.05. Bold font indicates ranges in the first 300ms of the P300.
Freq. Band
Range
Binding p-value
Binding Effect Size
Familial Theta
0.3-0.4, 0.4-0.5
0.028, 0.058
1.60, -1.12
Sporadic Delta
0.8-0.9
0.035
-0.55
Sporadic Theta
0.5-0.6
0.013
-1.19
Sporadic Alpha
0.1-0.2, 0.2-0.3, 0.5-0.6
0.09, 0.03, 0.03
-0.88, -1.17, -1.10
@@
Table 2 shows results that are of clinical interest after applying FDR correction at the 5 and 10 percent level. The main results that survive FDR correction are 0.3-0.6 range results in the lower frequency bands. There are overlapping significant results in the 0.3-0.6s range in the familial and sporadic AD data sets. The 0.3-0.6 results in the Sporadic AD alpha band pass FDR correction, thus bringing evidence of an ageing interplay between the alpha band frequency in the older sporadic AD patient group mimicking the behaviour of the Theta band in the younger familial AD patient group. The binding task effect sizes are all consistently greater in the patient group suggesting increased FAST connectivity in Alzheimer’s patients during binding VSTM tasks. This correlates to our time series plot of the FAST filtered mean patient and control matrices. The 0.3-0.4 range in the familial theta band shows greater squared difference values in controls with a rapid switch to greater values in patients in the next time step. We conjecture that this could be due to the oscillating behaviour of Event Related Potentials.
IV Discussion
@@
After applying these results to our two independent sporadic AD and Familial AD data sets, we found consistent overlapping clinically significant results in the 0.3-0.6 second range. This corresponds to the P300 range previously implicated in Alzheimer’s Disease [19]. Exploiting the binding deficit proposed by Pietto et al.[9] we have found evidence of an increase in anti-correlative activity or signal variation in Alzheimer’s patients in the P300 range during binding VSTM tasks and not in Shape VSTM tasks. This supports the binding task deficit as a potential biomarker for AD. FAST Connectivity performed well in negating spurious connections and emphasizing the important ones in the P300 range. Theta band irregularities have been well-researched to be linked to MCI due to Alzheimer’s disease[23]. While the slowing of the alpha band is an indicator of progressing AD [24]. The Alpha and Theta band ’shifts’ could be of clinical significance as the two independent datasets differ by age; thus the alpha band in the older Sporadic AD data set mimicking the behaviour of the Theta band in the younger Familial AD data set in the P300 range could provide an indicator for the progression of the disease from MCI to more severe full-on dementia. Alternatively, this frequency shift may be signaling age-related compensatory neural mechanisms which have been previously reported during memory tasks performed in the fMRI scanner [25].
Given our low sample size, we acknowledge the effectiveness of FAST Connectivity at picking out clinically significant results, especially given we have more controls in the sporadic AD data set. After rigorously performing simulations showing the effectiveness of our novel methodology, our application of it to the two data sets has provided potential biomarkers for the early detection of Alzheimer’s Disease and the progression of MCI to Dementia. Given the non-invasive nature of EEG signal Analysis combined with the low computational cost of using GVD connectivity with a relatively small number of patients, we see the potential for this method to be used in the diagnosis of Alzheimer’s disease for low-income individuals.We also see potential for accurate ERP detection at a lower number of trials, which could address the cognitive fatigue of participants in these working memory-based tasks. This could also reduce the cost, especially in low-income countries such as Latin America.
@@
An obvious application of FAST connectivity or similar methods would be in Brain Computer Interfaces (BCI).The ability to exploit discriminating information in real time from cheap, non-invasive EEG signals provides an avenue for an realistic, widely accessible BCI. While we are still far away from real-time detection, FAST connectivity, with its high in-variance to window length changes and performance at temporal resolution, shows potential for this one day being a possibility. Given recent advances in network based BCIs [26] and the proven importance of functional connectivity dynamics in the performance of BCIs [27] this would be a worthwhile avenue to explore.
Given the provision of more data, we could have drawn more validity to our clinical results. However they are still promising. Further work should focus on the optimization of the bi-variate node function and the Long-term Stable connectivity filter, focusing on interpretability and robustness against noise. In particular, it would be beneficial to look into the effect of certain node functions emphasizing specific signal properties such as anti-correlative information seen with the squared difference.The diagnosis of neurological disorders in psychiatry is an area of uncertainty due to the overlap between disorders. GVD Connectivity provides potential in the analysis of EEG signals to provide a more quantitative judgement on the nature of the neurological disorder. Machine Learning can be implemented on the network metrics of GVD connectivity due to the high temporal resolution of the metrics, this could add important transient information to Machine learning algorithms significantly improving performance akin to the wavelet transform shown to increase classification accuracy of neurological disorders by adding transient information[28].
@@
V Conclusion
We have introduced FAST connectivity, a novel algorithm building on Graph Variate Dynamic (GVD) Connectivity that leverages a single global filter computed from both groups of participants in a given EEG paradigm. We have shown, in controlled simulations on synthetic data, that the method outperforms previous Graph-Variate methods in detecting subtle differences in small temporal windows between groups of participants with and without a simulated ERP in noisy conditions. We have also shown the lower dependence on window length of the method, providing an alternative to existing sliding window methods. Of notable interest is the fact there is still relatively good performance when the window length is equal to the sampling rate allowing us to potentially detect changes in temporal windows at a very granular level. Applying FAST connectivity to two independent cohorts of Sporadic and Familial Alzheimer’s Disease patients engaged in VSTM tasks, we found significant differences between groups in the 0.3-0.6 second range in the binding task but not in the shape task, previous studies corresponds this to the P300 ERP range . This was more prominent in the lower frequency bands, in particular, the theta band. Corresponding with previous studies on the binding deficit and the role of the theta band in Alzheimer’s disease . Future work should aim to focus on further studying the properties of the Global filter, such as the Graph Laplacian eigenvalues in order to analytically understand its noise reduction and important connection promotion effects. Different node functions based on the structure of the graph signal data should also be explored.
VI Acknowledgements
Om Roy (Primary Author) is supported by the Engineering and Physical Sciences Research Council (EPSRC) Student Excellence Award (SEA) Studentship provided by the United Kingdom Research and Innovation (UKRI) council.
IA is partially supported by grants from ANID/FONDECYT Regular (1210195 and 1210176 and 1220995); ANID/FONDAP/15150012; ANID/PIA/ANILLOS ACT210096; ANID/FONDAP 15150012; and the MULTI-PARTNER CONSORTIUM TO EXPAND DEMENTIA RESEARCH IN LATIN AMERICA [ReDLat, supported by Fogarty International Center (FIC) and National Institutes of Health, National Institutes of Aging (R01 AG057234, R01 AG075775, R01 AG21051, CARDS-NIH), Alzheimer’s Association (SG-20-725707), Rainwater Charitable foundation – Tau Consortium, the Bluefield Project to Cure Frontotemporal Dementia, and Global Brain Health Institute)]. The contents of this publication are solely the authors’ responsibility and do not represent the official views of these institutions. The funders had no role in study design, data collection and analysis, publication decisions, or manuscript preparation.

### v2

@@ 3.2 Application to visual short-term memory binding in Alzheimer’s disease
EEG micro-states are transient patterns of the electroencephalogram that occur in very small temporal windows and are considered to be related to the most basic of human neurological processes. They have been previously shown to be able to distinguish between neurological disorders such as schizophrenia based on these tiny temporal window differences where the overall functional connectivity of the brain may be very similar [37]. Recently, there has been significant interest in these EEG micro-states in neurological disorder diagnosis.
@@
We ran experiments for the MCI-FAM and MCI-SPO datasets separately with a single FAST filter computed from participants in both the shape and binding tasks for each frequency band. As mentioned in our statistical methods section we computed the mean clustering coefficient and edge weights of each participant at ten disjoint temporal windows of the total one second epoch of interest and undertook non-parametric statistical testing between controls and patients to look for temporal windows where there is concurrently a significant difference between controls and patients in the binding task and no significant difference in the shape task. This exploits the proposed binding deficit established in [22]. Figure 4 shows the plots of the log of the $p$-values of the patients versus controls in shape and binding tasks for the MCI-FAM dataset with values below the red line indicating a significant difference.
@@
Figure 4: P-value plot for Controls vs Patients using FAST Connectivity in the Shape and Binding task in Delta(top),Theta(middle) and Alpha(bottom) bands for MCI-FAM
@@
The first thing we notice is that the behaviour of the delta and theta band in the binding task is nearly identical with significant results found at 0.3-0.4 seconds by the mean edge weights and 0.4-0.5 by the mean weighted clustering coefficient. We can see how having two different network metrics can aid the detection of clinically significant results. We see the shape tasks for these time steps are not significantly different thus these can be considered results of clinical interest as the specificity of binding deficits observed behaviorally are replicated here at a neural level. The binding task in the alpha band seems to show some behaviour similar to the theta and delta bands with a clinically interesting result at 0.4-0.5 seconds, however this would not pass FDR correction. The beta band seems to follow the same pattern in the 0.3-0.6 range with a dip towards the significance line in the binding task and movement away from it in the shape task. In light of volume conduction effects, the Gamma band was found to yield spurious results, consequently warranting its exclusion from our analysis. Overall, we notice a trend of clinically significant results in the 0.3-0.6 second range in the MCI-FAM data-set, mainly in the lower frequency bands.
@@
Figure 5: P-value plot for Controls vs Patients using FAST Connectivity in the Shape and Binding task in Delta(top),Theta(middle) and Alpha(bottom) bands for MCI-SPO
Figure 5 shows the plots for the Log of the p-values for patients vs controls in the shape and binding VSTM task for the MCI-SPO dataset against the time intervals.
@@
We notice similar patterns in the lower frequency bands with the binding task in the delta and theta band following similar patterns in the mean edge weights and the network metrics picking up clinically interesting results in the theta band at 0.4-0.5 and 0.5-0.6 seconds. This is an overlap in the 0.3-0.6 second range with the MCI-FAM data set. The delta band has a clinically interesting result at 0.5-0.6 seconds with the mean edge weights and a highly significant result at 0.8-0.9 seconds using the weighted clustering coefficient,this could be related to the emotion related Late Positive Potential (LPP). The LPP, characterized by a gradual positive shift in activation, typically manifests approximately 400-1000 ms following stimulus presentation. Its amplification has been linked to memory encoding and retention mechanisms [22]. Furthermore, it has been correlated with post-retrieval phases, such as decisional and evaluation processes which could be affected by AD.Again the overlap with the delta band in the MCI-FAM group in the 0.3-0.6s range is seen. Moreover,the alpha band has clinically interesting results at 0.1-0.2 seconds with the weighted clustering coefficient and 0.2-0.3s and 0.5-0.6s with the mean edge weights. It seems the similarity to the theta band in the MCI-FAM group is growing with the behaviour in the 0.3-0.6s range being much more prominent in the alpha band. The beta band has a significant result at the 0.4-0.5s time-step which is similar to the behaviour of the beta band in the MIC-FAM data-set (dip towards significance line in binding task, movement away in the shape task). Overall, there are consistent overlapping results of clinical interest in the 0.3-0.6s temporal range.
@@
We then applied FDR correction at the 10 percent and 5 percent level to account for multiple comparisons. Table 3 shows the time intervals at which there is a significant difference between controls and patients in the binding task and not in the shape task after applying FDR correction at the 5 and 10 percent level.
@@
Table 3: FDR Corrected Data Sets (10 percent level). The underlined text represents the Binding p-value below 0.05. Bold font indicates ranges in the first 300ms of the P300.
Freq. Band Range Binding p-value Binding Effect Size
MCI-FAM Theta 0.3-0.4, 0.4-0.5 0.028, 0.058 1.60, -1.12
MCI-SPO Delta 0.8-0.9 0.035 -0.55
MCI-SPO Theta 0.5-0.6 0.013 -1.19
MCI-SPO Alpha 0.1-0.2, 0.2-0.3, 0.5-0.6 0.09, 0.03, 0.03 -0.88, -1.17, -1.10
@@
The main results that survive FDR correction are 0.3-0.6 range results in the lower frequency bands. There are overlapping significant results in the 0.3-0.6s range in the MCI-FAM and MCI-SPO data sets. The 0.3-0.6 results in the MCI-SPO alpha band pass FDR correction, thus bringing evidence of an ageing interplay between the alpha band frequency in the older MCI-SPO patient group mimicking the behaviour of the Theta band in the younger MCI-FAM patient group. The binding task effect sizes are all consistently greater in the patient group suggesting increased FAST connectivity in Alzheimer’s patients during binding VSTM tasks. This correlates to our time series plot of the FAST filtered mean patient and control matrices. The 0.3-0.4 range in the MCI-FAM theta band shows greater squared difference values in controls with a rapid switch to greater values in patients in the next time step. We conjecture that this could be due to a delay in the onset of the P300 in patients resulting in the increased FAST connectivity to only appear after the onset of the P300 in the controls in the previous time-step.
Furthermore, we computed the average of the FAST filtered connectivity tensors for all participants and controls in the binding task for both MCI-FAD and MCI-SPO. This gave us one tensor for patients and one for control representing general activity in terms of functional connectivity in the task.
We mapped the node-wise connectivity to specific electrodes and compared controls and patients during a specific time-step in the P300 range, in particular we set the the time-step corresponding to around the 0.5 second mark in the Theta band in both datasets as this was the overlapping time-step that was common to both datasets in the Theta frequency band with similar effect sizes.
(a) MCI-FAM
(b) MCI-SPO
Figure 6: Top 1 Percent (MCI-FAM) and Top 0.1 percent (MCI-SPO) FAST Connectivity Topological Scalp Plots in the Theta Band (Binding Task) at 0.5 seconds Post-Stimuli Onset (Intensity Indicated on Colour bar)
We noted that only the top 1 percent of connections in the union of both mean matrices for the binding task appeared in both Controls and Patients in MCI-FAD, while in MCI-SPO there were similarly strong connections up to the top 0.1 percent of connections. This is due to both controls and patients in MCI-SPO having relatively large values for the instantaneous local variation with patients having more connections. In MCI-FAM,on the other hand,patients had typically greater local variation and more connections with a larger local Dirichlet Energy compared to controls.
Figure 6 thus shows the connectivity plots where we compute the top 1 percent and 0.1 percent of strongest connections in the MCI-FAM and MCI-SPO mean matrices respectively.
Interestingly we notice an increase in instantaneous variability in patients in both MCI-FAD and MCI-SPO compared to controls in the P300 response. In particular we can empirically observe the increased presence of strong potential ’hubs’ where multiple strongest connections start from the same node.
In MCI-FAD these hubs originate from the PO3 and PO6 electrodes and is strongly connected to fronto-central regions. There is also a hub from FC3 to the parietal-occipital region. In general we notice increased local variation from the parietal-occipatal to the fronto-central region.In MCI-SPO there is clear major hub originating at C1 with large local variation with regions across the entire parietal area, D1 is similar to C1 yet slightly less prominent and smaller hubs also originate from across the parietal region. We notice the difference between MCI-SPO and MCI-FAD being the increased Local Dirichlet energy is more widespread across the parietal area.
These empirical results are fascinating but is not the focus of our study where we are more focused on global connectivity changes. The thorough broad spatial origins of these functional connectivity alterations however seems to be a promising avenue for future research.
4 Discussion
@@
After applying these results to our two independent MCI-SPO and MCI-FAM data sets, we found consistent overlapping clinically significant results in the 0.3-0.6 second range. This corresponds to the P300 range previously implicated in Alzheimer’s Disease [33, 34].We also found evidence of altered functional connectivity related to increased localized signal variation in Alzheimer’s patients in the P300 range at time-steps specific to the binding task. This supports the binding task deficit as a potential biomarker for AD. Theta and Alpha band irregularities have been well-researched to be linked to cognitive dysfunction and MCI due to Alzheimer’s disease[40, 44, 43]. While the slowing of the alpha band is an indicator of progressing AD [41]. The Alpha and Theta band ’shifts’ could be of clinical significance as the two independent datasets differ by age; this frequency shift may be signaling age-related compensatory neural mechanisms which have been previously reported during memory tasks performed in the fMRI scanner[42]. More importantly, we note that our results correspond to what is observed at clinical performance level of the VSTM tasks with binding task performance being significantly worse in patients and controls in both MCI-SPO and MCI-FAD. We have thus confirmed at a neuronal level what we see in practice.
It is of interest to note that the increased local Dirichlet energy from parietal to central-frontal regions is consistent in both types of AD compared to controls and these are distinctly related to functional connectivity between non-structurally related changes indication functional connectivity changes could indicate AD progression long before structural changes occur. These hubs could indicate neuronal level abnormalities of excitation and inhibition that are shown to be associated with tau and amyloid beta in pre clinical models of AD[38] .
It is promising that our simulations used a lower electrode density (where we employed 31 electrodes) and FAST connectivity was still able to reliably pick up simulated ERP’s at a relatively low number of trials, this brings further confidence to our AD data results that employ 64 and 128 electrodes for the MCI-FAM and MCI-SPO data sets respectively, indicating that the electrode density had a fairly negligible influence on our results, again this corresponds to [29] that stated that performance on functional connectivity measures was negligible above 60 electrodes.
While FAST connectivity is not in any way meant to replace power based analysis of ERP’s it’s significantly improved performance in this specific time-locked task simulation shows that it can be a reliable tool to analyze dynamic functional connectivity changes related to ERP’s at very small time scales.
The data we have utilized in this work is a small sample of two populations with different risk levels (Mild Cognitive Impairment) for AD (E280A-PSEN1 Familial AD, with 100 percent risk [47, 50] and sporadic MCI with an unknown risk. Previous work has shown that EEG connectivity can distinguish mutation carriers from controls with accuracy near 90 percent [53] a classification accuracy never reached via pure behavioural scores [51]. It has also been shown that the EEG features linked to Visual Short-Term Memory Binding deficits of these patients across the two variants are indistinguishable [51]. We have shown that FAST connectivity can distinguish the subtle time-frequency changes between MCI-SPO and MCI-FAM. Parra et al.[52] also recently showed that a cost of binding (drop in performance on the shape-colour binding condition relative to memory for shapes only) greater that 20 percent was associated to increased Amyloid Beta deposits in still cognitively unimpaired older adults.
In light of this, our application of FAST connectivity to the MCI-SPO and MCI-FAM datasets have provided results of interest for understanding deficits of VSTM binding in AD. With more data and further analysis (including at an individual,rather than just group level),this could potentially also be useful as a diagnostic indicator for the early detection of Alzheimer’s Disease and the progression of MCI to Dementia. Given the non-invasive nature of EEG signal Analysis combined with the low computational cost of using GVD connectivity with a relatively small number of patients, we see the potential for this method to be used in the diagnosis of Alzheimer’s disease for low-income individuals. Furthermore, The task has been recommended by international consensus groups [48] as a promising pre-clinical test for Alzheimer’s disease. Some have already introduced the task in their clinical practice The task has now been introduced in major international cohort studies such as PREVENT [54] and RedLAT [49].
Working memory tasks often require participants to engage in sustained cognitive effort, leading to potential cognitive fatigue, especially in prolonged experimental sessions. FAST connectivity shows potential to address this challenge by exploring the feasibility of achieving accurate ERP detection with a lower number of trials. Additionally, we should take into account economic considerations prevalent in low-income countries, where optimizing experimental protocols can significantly reduce costs associated with data acquisition and analysis.
@@
An obvious application of FAST connectivity or similar methods would be in Brain Computer Interfaces (BCI) [46].The ability to exploit discriminating information in real time from cheap, non-invasive EEG signals provides an avenue for an realistic, widely accessible BCI. While we are still far away from real-time detection, FAST connectivity, with its high in-variance to window length changes and performance at temporal resolution, shows potential for this one day being a possibility. Given recent advances in network based BCIs [45] and the proven importance of functional connectivity dynamics in the performance of BCIs [46] this would be a worthwhile avenue to explore.
Machine Learning can be implemented on the network metrics of GVD connectivity due to the high temporal resolution of the metrics, this could add important transient information to Machine learning algorithms significantly improving performance akin to the wavelet transform shown to increase classification accuracy of neurological disorders by adding transient information[36].[39] also showed that combining graph metrics based on dynamic functional connectivity in small temporal windows with typical classification algorithms showed significantly improve performance in the early detection of Parkinson’s Disease (PD), showing the discriminating ability of these EEG micro-states.
The diagnosis of neurological disorders in psychiatry is an area of uncertainty due to the overlap between disorders. FAST Connectivity provides potential in the analysis of EEG signals to provide a more quantitative judgement on the nature of the neurological disorder.
@@
5 Conclusion
We have introduced FAST connectivity, an algorithm that leverages a single global filter computed from both groups of participants in a given EEG paradigm. We have shown, in controlled simulations on synthetic data, that the method outperforms previous Graph-Variate and traditional power spectra methods in detecting subtle differences in small temporal windows between groups of participants with and without a simulated ERP in noisy conditions. We have also shown the lower dependence on window length of the method, providing an alternative to existing sliding window methods. Of notable interest is the fact there is still relatively good performance when the window length is equal to the sampling rate allowing us to potentially detect changes in temporal windows at a very granular time-scale while also showing potential to reduce the required number of trials required for ERP analysis. Applying FAST connectivity to two independent cohorts of Sporadic and Familial MCI patients engaged in VSTM tasks, we found significant differences between groups for time steps in the 0.3-0.6 second range in the binding task but not in the shape task, previous studies corresponds this to the P300 ERP range.This was more prominent in the lower frequency bands, in particular, the theta band. Corresponding with previous studies on the binding deficit and the role of the theta band in Alzheimer’s disease and general dysfunction in memory and cognition. Future work should aim to focus on further studying the spectral properties of the FAST filter, such as the spectral profile of the Graph Laplacian in order to analytically understand its noise reduction and important connection promotion effects. Different instantaneous node functions based on the structure of the graph signal data should also be explored.
Acknowledgements.
Om Roy (Primary Author) is supported by the Engineering and Physical Sciences Research Council (EPSRC) Student Excellence Award (SEA) Studentship provided by the United Kingdom Research and Innovation (UKRI) council. IA is partially supported by grants from ANID/FONDECYT Regular (1210195 and 1210176 and 1220995); ANID/FONDAP/15150012; ANID/PIA/ANILLOS ACT210096; ANID/FONDAP 15150012; and the MULTI-PARTNER CONSORTIUM TO EXPAND DEMENTIA RESEARCH IN LATIN AMERICA [ReDLat, supported by Fogarty International Center (FIC) and National Institutes of Health, National Institutes of Aging (R01 AG057234, R01 AG075775, R01 AG21051, CARDS-NIH), Alzheimer’s Association (SG-20-725707), Rainwater Charitable foundation – Tau Consortium, the Bluefield Project to Cure Frontotemporal Dementia, and Global Brain Health Institute)]. The contents of this publication are solely the authors’ responsibility and do not represent the official views of these institutions. The funders had no role in study design, data collection and analysis, publication decisions, or manuscript preparation.